Given a standard normal distribution with mean =118 and standard deviation = 29, find P(76 < X < 144)


1
Expert's answer
2022-01-25T14:50:08-0500

Given "\mu=118, \sigma=29" then,

"p(76\lt x\lt 144)=p({76-118\over29}\lt Z\lt{144-118\over29})=p(-1.45\lt Z\lt 0.90)"

This is equivalent to,

"p(-1.45\lt Z\lt 0.90)=\phi(0.90)-\phi(-1.45)=0.8159-0.0735=0.7424"

Therefore,

"p(76\lt x\lt 144)=0.7424"